Backlinks to ArmandoRiccardi in all Webs (Search Main Web only)

No backlinks in this web

Results from AdaptiveOptics web retrieved at 10:23 (GMT)

2011
2011 * Minutes_from_LBT672_progress_meeting_14_Jan_2011.pdf: Meeting Minutes 2011 01 14 * Minutes_from_LBT672_progress_meeting_3_Feb_2011.pdf: Meeting Minutes...
Number of topics: 1

Results from FLAO web retrieved at 10:23 (GMT)

AcqInfFunc4dAndFlattening
Acquisition of influence functions and flattening Acquisition of influence functions Acquisition of raw interferograms Preliminary operations: 1 Verify that ...
AdOptMeetings
Progress and Revision Meetings * LBT AO Progress Meeting. Arcetri 21 24 Feb. 2005 * LBT AO Review Meeting. Arcetri 10 11 Nov. 2005 * LBTO Software Worksh...
AdSecLogBrowsing
AdSec Log Bowsing 1 cd $ADOPT_LOG 1 less FASTDGN00.log 1 search for EMERGENCY (first of the sequence around the time in which the problem happened) 1 ...
AdsecCommissioning
LBT672a Commisioning LogBook day by day Pre summer shutdown runs LabLog672a20100220 LabLog672a20100221 LabLog672a20100222 LabLog672a20100223 LabLog672a20100224 ...
AdsecCommissioning672b
Unit B: System integration in Clean Room July 2011 Run: system integration in Clean Room * CleanroomIntegration672b20110701 * CleanroomIntegration672b20110...
CheckFirmwareDisablingCoils
How check the firmware safety functions (distance threshold and CrateBCU watchdog) Distance thresholding for testing the counter of the occurrences of distance t...
CleanRoomIntegration672a20110819
Unit B: 19 Aug 2011 activities * Present people: * Arcetri: Armando Riccardi, Marco Xompero * LBTO: Guido Brusa, Doug Miller, Michael Wagner, Richa...
CleanRoomIntegration672a20110820
Unit B: 20 Aug 2011 activity * cleaning magnets and shell (shell on the stand in the clean room) * shell blown with dry Nitrogen * cleaning top sur...
CleanRoomTest672b20111113
* conf/adsec/672b/configuration.txt: all null string except shell field. Old file saved in configuration.txt.copy * conf/adsec/672b/TS5/elec.txt: empty act_wo_...
CleanRoomTest672b20111118
Procedure: * remember to people on the summit to: * switch on the main cooling * switch on the Neslab * check 110V is already plugged in ...
CleanRoomTest672b20120624
modi forzati: 647 reference flattening con la vecchia scheda: /local/aomeas/adsec_calib/status_save/20120709_135656/status_save_flat.sav Flattening con la nuova s...
ConFiguraTions
Configurations FLAO#2 FLAO#2 @ Left Rear Bent Gregorian M3 position Official position, confirmed 2012 02 Tip Tilt Piston Sel Date Note 1115 ...
FLAO1TelLog20141101
01 Nov 2014. Daytime activity RR montato. AGW montato da ieri. Allineamento con il RR Rimosse protezioni dal collimation mirror e dalla flowerpot. Power on W: ...
FLAO1TelLog20141102
h7:30 Retro ON Abbiamo homato e riposizionato esapodo e terziario ma lo spot non torna nel W. AO print, hexa_spiral(30,15,/apply) Andiamo a controlare le viti de...
r9 - 29 Sep 2015 - 14:03 by EnricoPinna
FLAO2LabLog20120329
Daytime#1 Dayrun#1. Lorenzo, Carmelo, Fernando Armando h11:00 Power On AdSec Adsec era gia' on. Settata la shell, posizionato M3 e HP nelle posizioni definite i...
r4 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120331
Daytime#3 Dayrun#3. Lorenzo, Carmelo, Fernando, Armando Inizio operazioni: h08:15 ADSEC Rimossi act 20 e 582 perche' causavano RIP ieri sera alla fine della gior...
r9 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120401
Daytime#4 Dayrun#4. Lorenzo, Carmelo, Fernando, Armando partenza alle h8:15 BIN1 per controllare che la densita' installata ieri non dia problemi e per rifare i...
r7 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120402
Daytime#5 Dayrun#5. Lorenzo, Carmelo, Fernando, Armando Allineamento esapodo di oggi x y z Rx Ry Rz 4.4 1.7 4.45 306 592 0 Il rirotato...
r7 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120403
Nighttime#1 Dayrun#6. Lorenzo, Carmelo, Fernando, Armando 02:50UT 04/Apr/UT (19:50 03/Apr local time) Il telescopio stenta a partire. DIMM non funzionante, ma, s...
r8 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120404
Nighttime#2 Dayrun#7. Lorenzo, Carmelo, Fernando, Armando 02:20UT 05/Apr (19:20 04/Apr Local time) UTH both OFF. M1 ventilation OFF AGW cover ON AO366 Modificat...
r11 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120405
Nighttime#3 Dayrun#8. Lorenzo, Carmelo, Fernando, Armando E' stata cambiata la versione del TCS al BP14, quella usata a febbraio in cui funziona l'offset. Il roll...
r8 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120406
Nighttime#5 Dayrun#8. Lorenzo, Carmelo, Fernando, Armando 02:30UT 07/Apr/UT (19:30 06/Apr local time) Cielo velato. Potrebbe essere un problema per testare il fai...
r6 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120407
Nighttime#6 Dayrun#9. Lorenzo, Carmelo, Fernando, Armando 02:15UT 08/Apr/UT (19:15 07/Apr local time) Cielo velato. ND rimossa: * riconfigurato lo zero point ...
r4 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120408
Daytime#6 Dayrun#10. Lorenzo, Carmelo, Fernando, Armando BUG: AdSec. Quando viene dato un Correct MODES (es. astigmatismo), se la LUT dell'astigmatismo e' attiava...
r4 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2LabLog20120916
Night time: UT 2012 Sep 16 Ora attesa per iniziare: 04UT (21LT, prima Doug lavora su nostro AGW) Sorgenti selezionate: * AO741: culmina alle 5:30UT ( 70deg el...
FLAO2LabLog20120917
2012 Sep 17 UT Simone, Lorenzo, Armando Modifica al file di slop null. Dovrebbe eriolvere il problema delle slope rms a Inf tree house OFF mirror ventilation ON ...
FLAO2LabLog20120918
18 Sep 2012 UT Simone, Lorenzo, Armando, Juan Carlos Abbiamo tentato di usare la nuova versione di Alfio del software (adsec e wfs), ma all'adsc_start abbiamo avu...
FLAO2LabLog20120919
2012 Sep 19 UT Simone, Lorenzo, Armando, Juan Carlos 03:00UT DIMM 1.0 AO734 03:00UT (mag 15.3 con ND) 03:18UT DIMM 0.8 Siamo ancora sulla versione del oftware ...
FLAO2NightProcedures
Proc: inizio notte 1. Controllare RR nel file di conf 1. Controllare se la ND e' intallata. 1. Controllare il settaggio dello zero point della magnitudin...
r5 - 11 Sep 2012 - 16:48 by EnricoPinna
FLAO2Run3ObsPlan
Programma per la notte #3 * BIN1, 500 modi, 990Hz Programma della prima notte (miseramente fallito a causa del Guasto della BCU crate#1) * BIN1 400 modi ta...
r5 - 11 Sep 2012 - 16:48 by EnricoPinna
Flao2AcceptanceActionItems
FLAO#2 Acceptance test Action Items Priority levels * 0: before dismounting the FLAO#2 system from the tower (currently scheduled on 23/May) * 1: before st...
Flao2AcceptanceDocuments
FLAO#2 Acceptance documentation Main.ArmandoRiccardi 17 May 2011 * SW_Acceptance_Test_Unit_2_640s013b.pdf: LBT Adaptive Optics System Acceptance Test. Unit...
FlaoActivityAtLbtoReports
General FLAO activity daily reports * FlaoActivityReport20100215 * FlaoActivityReport20100216 * FlaoActivityReport20100217 * FlaoActivityReport2010021...
FlaoActivityReport20100215
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO whe...
FlaoActivityReport20100216
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100217
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO we...
FlaoActivityReport20100218
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100219
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100220
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100221
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100222
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100225
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100226
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport20100227
Dear Richard dear all, please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO wer...
FlaoActivityReport201002280301
Dear Richard dear all, please find below a short summary of the activities done Sunday today on the mountain related to the AO system. Italian people working at...
FlaoActivityReport20100302
Dear Richard dear all, please find below a short summary of the activities done Sunday today on the mountain related to the AO system. Italian people working at...
FlaoActivityReport20100303
Dear Richard dear all, please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were: ...
FlaoActivityReport20100304
Dear Richard dear all, please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were: Ar...
FlaoActivityReport20100305
Dear Richard dear all, please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were: Ar...
FlaoActivityReport20100306
Dear Richard dear all, please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were: Ar...
FlaoActivityReport20100307
Dear Richard dear all, please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were: Ar...
FlaoActivityReport20100309
Dear Richard dear all, please find below a short summary of the activities of today on the mountain related to the AO system. Italian people working at LBTO were:...
FlaoActivityReport20100310
Dear Richard dear all, please find below a short summary of the activities of Sunday Monday on the mountain related to the AO system. Italian people working at ...
FlaoActivityReport20100311
Dear Richard dear all, please find below a short summary of the activities of today on the mountain related to the AO system. Italian people working at LBTO were:...
FlaoActivityReport20100312
Dear Richard dear all, please find below a short summary of the activities of today on the mountain related to the AO system. Italian people working at LBTO were:...
FlaoActivityReport2010031314
Dear Richard dear all, please find below a short summary of the activities of 13th and 14th on the mountain related to the AO system. Italian people working at L...
FlaoActivityReport20100315
Dear Richard dear all, please find below a short summary of the activities of 15th on the mountain related to the AO system. Italian people working at LBTO were:...
FlaoActivityReport20100316
Dear all, please find below a short summary of the activities of march 16th on the mountain related to the AO system. March 16th Italian people working at LBTO we...
FlaoActivityReport20100317
Dear all, please find below a short summary of the activities of 17th on the mountain related to the AO system. Italian people working at LBTO were:L.Fini, M.Xom...
FlaoActivityReport20100318
Dear all, please find below a short summary of the activities of 17th on the mountain related to the AO system. Italian people working at LBTO were:L.Fini, M.Xom...
FlaoActivityReport20100319
Dear all, please find below a short summary of the activities of 19th on the mountain related to the AO system. Italian people working at LBTO were:L.Fini, M.Xom...
FlaoActivityReport20100320
Dear all, please find below a short summary of the activities of 19th on the mountain related to the AO system. Italian people working at LBTO were:L.Fini, M.Xom...
FlaoActivityReport20100321
Dear all, please find below a short summary of the activities of 21th on the mountain related to the AO system. Italian people working at LBTO were:L.Fini, M.Xom...
FlaoActivityReport20100322
Dear all, please find below a short summary of the activities of 19th on the mountain related to the AO system. Italian people working at LBTO were: M.Xompero ...
FlaoActivityReport20100323
Dear all, please find below a short summary of the activities of 23th on the mountain related to the AO system. Italian people working at LBTO were: M.Xompero ...
FlaoActivityReport20100324
Dear all, please find below a short summary of the activities of 24th on the mountain related to the AO system. Italian people working at LBTO were: M.Xompero,...
FlaoActivityReport20100325
Dear all, please find below a short summary of the activities of 25th on the mountain related to the AO system. People working at LBTO: G. Brusa people working...
FlaoActivityReport20100326
Dear all, please find below a short summary of the activities of the 26th related to the AO system. People at LBTO: G. Brusa, D. Miller people working in remote...
FlaoActivityReport20100327
Dear all, please find below a short summary of the activities of the 27th related to the AO sys People at LBTO:G. Brusa people working in remote:A. Riccardi (V...
FlaoActivityReport20100328
Dear all, please find below a short summary of the activities of the 27th related to the AO system.People at LBTO:G. Brusa people working from remote:A. Riccard...
FlaoActivityReport20100329
Dear all, please find below a short summary of the activities of the 28th related to the AO system. People at LBTO:G. Brusa people working from remote:A. Ricc...
FlaoActivityReport20100330
Dear all, please find below a short summary of the activities of the th related to the AO system. People at LBTO: R. Briguglio reote support A. Riccardi an...
FlaoActivityReport20100331
Dear all, please find below a short summary of the activities of the the related to the AO system. People at LBTO: R. Briguglio remote support A. Riccardi ...
FlaoActivityReport20100401
Dear all, please find below a short summary of the activities of the th related to the AO system. People at LBTO: R. Briguglio remote support M.Xompero, A....
FlaoActivityReport20100403
Dear all, please find below a short summary of the activities related to the AO system. People at LBTO: R. Briguglio remote support M.Xompero, E.Pinna(VNC ...
FlaoActivityReport20100404
Dear all, please find below a short summary of the activities related to the AO system. People at LBTO: R. Briguglio remote support A.Riccardi (VNC Skype) ...
FlaoActivityReport20100405
Dear all, please find below a short summary of the activities related to the AO system. People at LBTO: R. Briguglio Summary Operations began at 1:40UT, w...
HowToAdSecMirror
HOW TO How to: * setup variables by hand for wind and elevation: * create variables public/float R.AOSR00.AMB.WINDSPEED (values in m/s, 8 is the TSS thre...
HowToLBT672
HowTo for LBT672 units * Lbt672SoftwareOnWin: How to run the LBT672a software on Windows * HowToSetupLbt672InCleanRoom: How to setup and test the mirror uni...
HowToSetupLbt672InCleanRoom
How to setup and test the mirror unit in clean room 1 the mirror unit is supposed on the integration stand (see 642xnnna) and connected to the cooling system ...
LBT672aLogFiles
Log and Dump files Log Files At telescope log files for the dx mirror are in adsecdx workstation in directory $ADOPT_LOG * ADAMHOUSEKPR00: generated by xxx. D...
LBT672aSWModificationRequests
LBT672a requested software modifications Legenda 0 3: priorità (min=3, max=0) F = fatto Modification list 2 salvare il nome del file di feed forward nel file ...
LBT672aShippingList
Shipping List for the LBT672a Parti da rimontare sull'unità prima dell'uso e relativi tools * 3 settori retention ring con corrispondenti viti (12) ( Consegna...
LabLog20141029
Armando, Tommaso, Enrico, Mike L., Jonathan Sommario errori di allineanto W LUCI window * Alignm error table.png: Qui riportiamo gli erorri stiamti a questo...
LabLog20141030
h08:50 verifica allineamento e misura shims con Laser di Paolo Procedura per la shimmatura finale (appunti per la procedura da rilasciare a LBTO): 1 allentate ...
r5 - 01 Nov 2014 - 03:26 by EnricoPinna
LabLog672a20100227
27 Feb 2010 * Removed stiffening ring for the flipping * Joined siggen cables * verification RB to CP distance with dial gauge for excluding RB degluin:m...
LabLog672a20100313Night
Main.RunaBriguglio 13 Mar 2010 01:53 the system was already powered on load program setpistoncurrent(35) set_flat @30 deg elevation failed problema con TSS e abil...
LabLog672a20100915
Mounted miniskirt and removed shell mounting posts 11:00: start of electromechanical tests Tests with old configurations (calibrations, FF, etc.) * shell sett...
LabLog672a20100916
Noise of act 484. Investigation the visual inspections showed one wire of the coil touching the "orpello" inside the refplate hole. Possible source of interfere...
LabLog672a20100917
Problem with ADAM the problem found yesterday has been verified to be a problem in the adam_wrapper lib. SOLVED. It is still not understood because this criticali...
Lbt672AcceptanceDocumentRepository
Document repository of LBT672a/b Acceptance Phases The document and data repository related to the Acceptance phases of the LBT672a/b units is accessible via SFTP...
Lbt672ElecronicsSchemes
* Backplane v1.0 * Cables * Capacitive sensors (external) v1.0 * Capacitive sensors (internal) v2.0 * BCU Boards v1.0 * DSP Boards v2.0 * PCI Bo...
Lbt672Electronics
LBT672 Electronics Color codes reference signal wires Ref signal wires Crate# refsignal wireLBT672a refsignal wire MAG TS1 wire TS4 (TBC) TS3/TS5 wire ...
Lbt672Mechanics
LBT672 Mechanics General * ADS International s.r.l. "f/15 Adaptive secondary mechanical design" CAN#640a005f, 17 Dec 2003 * mirror_numbering_geometry.ppt: ...
Lbt672Optics
LBT672 Optics Drawings * Official repository: LBT CAN Reference Plate * Steve Miller (Mirror Lab) "LBT Reference body #3 Manufacturing Report" CAN#645s0...
Lbt672SchemesBackplane
Backplane Schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesBcu
BCU board schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesCables
Cable Schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesCapsensExternal
Capacitive sensors (External Boards) schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesCapsensInternal
Capacitive sensors (Internal Boards) schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesDsp
DSP board schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesPCI
PCI board schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesPowerBackplane
Power Backplane schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672SchemesReferenceSignal
Reference Signal Generator board schemes Main.ArmandoRiccardi 20 Mar 2007
Lbt672ShellWindProtectionMeetings
LBT672 Shell Protection Meetings * Meeting_minutes_Wind_protection.pdf: Minutes of the Wind Protection meeting of 23/12/2006 * Meeting_minutes_Wind_protecti...
Lbt672SoftwareBugs
Software Bugs and requested changes Housekeeper GUI 1. Summary System Tab: label asse x dei plot: minuti non in secondi 1. XXX Mirror GUI 1. XXX f...
Lbt672SoftwareOnWin
How to run LBT672 software on Windows 1 install in a directory ("mydir", herafter) the following directories from the SVN repository: idl, calib and conf 1 ...
Lbt672SparePartsInOAA
List of Spare components in OAA Last update : Main.ArmandoRiccardi 28 Jan 2010 ID Description Quantity Notes Additional Notes ELA098 Internal ...
Lbt672aGlycolAccidentReview
Review of Glycol Accident with LBT672a Reviewer Document. CAN#640a021a Main.ArmandoRiccardi 03 Apr 2009
LbtAoProgressMeeting03May2007
LBT AO progress meeting. Skype conference, 03 May 2007 * Minutes of the meeting: Main.ArmandoRiccardi 07 Jun 2007
LbtAoProgressMeeting04Jun2007
LBT AO progress meeting. Skype conference, 04 June 2007 * Minutes of the meeting: Main.ArmandoRiccardi 07 Jun 2007
LbtAoProgressMeeting11Jul2007
LBT AO progress meeting. Skype conference, 11 July 2007 * Minutes of the meeting: Main.ArmandoRiccardi 16 Jul 2007
LbtAoProgressMeeting30Mar2007
Monthly LBT AO progress meeting Skype conference on 30 March 2007 * Minutes of the meeting Main.ArmandoRiccardi 10 Apr 2007
LbtAoProgressMeetingAug2006
LBT AO progress meeting. Arcetri, 22 23 Aug. 2006 * Minutes of the meeting * LBT AO Updated schedule (Microsoft Project file) Main.ArmandoRiccardi 20 Ma...
LbtAoProgressMeetingMar2007
Minutes from LBT AO progress meeting Author: Joar Brynnel Date: March 6, 2007 Location: Arcetri I. Issues from last meeting CCD47 for W#2: The CCD47 #2 was not f...
LbtAoSchedules
LBT AO schedules Schedules are now included in every document of minutes of the montly LBT AO progress meetings. Go here and look for the most recent minutes docu...
LbtLab
Logs and notes of FLAO Commissioning . Work in clean room Lab log book of the work done in the clean room of the LBTO in Mt. Graham FLAO general reports (distrib...
LbtoMeetingExtractAug2006
Follow up of the August 2006 meeting All, This email is written as follow up on last week's discussions on the LBT672 schedule. This morning we had a meeting at t...
LogBook20090818
Ripristinato il salvataggio sulla SDRAM Il test ha lo scopo di verificare se il problema del blocco persiste nel caso in cui non vengano fatte letture dalla memor...
LogBook20090819
FLTimeout before starting test_time_hist2 Abbiamo avuto Il timeout di fibra prima di avviare test_time_hist2, dopo aver settato la shell con l'update_panels conti...
LogBook20090820
test_time_hist2 11:50 started test same logics as before. Storage of diagnostics in SDRAM enabled: IDL print, dspstorage() INCREASED the refresh frequency (i.e....
LogBook20090821
Test of new logics. Debug release "e" The new logics seems to corrupt the communication of the master diagnostics record No time for further investigation and/or...
LogBook20090901
System started at 11:20 with the following configuration: * Logics version: 2.35 (BCU) and 6.04e (DSP) * Same disturbance as before * Storage of diagnost...
LogBook20090902
11:55: started test with following configuration: * Logics version: 2.35 (BCU) and 6.04e (DSP) * Same disturbance as before * Storage of diagnostics in D...
LogBook20090903
11:42: started test with following configuration: * Logics version: 2.35 (BCU) and 6.04e (DSP) * Same disturbance as before * Storage of diagnostics in D...
LogBook20091014
Membrane Gluing at ADS Gluing procedure started (membrane insertion with glue) at 18:43: data saved with drift_test procedure. from 18:43 to 19:35 data set "a" in...
LogBook20091102
LBT672b in the hub * new routine for setting the TS4 gains: * print, ramp_gain_script60TS4() with no ICC and gains set as reported above, we measured: ...
LogBook20091103
Drift test a drift test has been launched for a nighttime test session. The housekeeper switched the unit off twice because of an overtemperature of the switch BC...
LogBook20091208
Simulation of Seeing Limited Acceptance Test routine for reducing data: $ADSEC_IDL_SOURCE/adsec_lib/adsec_utilities/acctest_sl_mode.pro flattening command file: ...
LogBook20101015B
from 10:30 to 11:30 the shell TS5 has been dismounted: * NO detached magnets shell re installed without miniskirt. central membrane screws tightened 160000 fsm...
MasterDiagnosticRecord
Master Diagnostic Frame The structure of the Tosetti's circular buffer that is dumped in case of failure All the definitions are in $ADOPT_SOURCE/lib/AdsecDiagnLi...
NotesCommissioning
Notes on procedures after experience at Commissioning Shell installation * add to unplug fiber, flux meter and hub temperature cables at the beginning * ad...
NotesForNextUnit
Notes for a better run with next unit 1 Clean Room: * Commit a corner for dressing (the best when entering on the right); hooks for suits, box for used ...
NumericalRecipesInC
Numerical recipes in C (2nd ed.) * Numerical_Recipes_in_C (collection of zipped PDF files) Main.ArmandoRiccardi 08 Oct 2008
PsfUseWithRrhHelp
Help on PSF use for alignemnt during AO calibration The example is written for the right side of the telescope (FALO#1), change right with left for the other side...
RoarksFormulasForStressAndStrain
Roark's Formulas for Stress and Strain (7th Edition) * 2542X_fm.pdf: Front Matter * 2542x_lot.pdf: List of Tables * 2542x_pref7.pdf: Preface to 7th edit...
SummaryMeeting20100412
Summary of meeting held in "Aula Rosa" on 12 Apr 2010 The discussion showed the following high priority tasks: * AO Arbitrator * fix the management of f...
TelLogLBT672a20100313
Time is UTC R.AOSR00.AMB.WINDSPEED has been set to 10.0 for enabling TSS =========== DX and SX shutters open Lateral and back vent closed External wind: about...
TelLogLBT672a20100325
Night Observation 25 26 Mar 2010 Local operator: Guido Brusa Remote operator: Armando Riccardi Main.ArmandoRiccardi 26 Mar 2010
TelLogLBT672a20100327
Primo reset della shell: FASTDGN00 INF 4699149 2010 03 28 05:08:32.309325 MAIN HighOrder offload averaging 12 FastDiagnostic.cpp:1199 FASTDGN00 INF 469915...
TelLogLBT672a20100328
Main.ArmandoRiccardi 29 Mar 2010
TelLogLBT672a20100330
Adaptive Secondary for RFBG Commissioning 30 March 2010 Adaptive Mirror: RBriguglio (INAF) Observer: JHeidt, JHill (LBTO)Telescope Operator: WWackSoftware: TEdg...
TelLogLBT672a20100404
Tests show that the RIP just after setting and flattening happen because the HO offload is enabled and applyed too short time after the enabling. The execution of...
TelLogLBT672a20100529
Mirror rip during closed loop FASTDGN00 ERR 738858 2010 05 29 04:00:32.482753 FUNCTEMERGENCYST CHCURRAVERAGE 0262 = 0.8 /home/aoacct/releases/src2/DiagnApp...
TelLogLBT672a20100601
Mounted RRH BentGregorianBack (interferometer focal station) New alignment required: Hexapod: x y z Rx Ry Rz mm mm mm as as as 1.3 ...
TelLogLBT672a20100623
2010 Jun 23, 8:20 UT Moving M3 toward interferometer focal station: M3 was previouslu configured for FLAO/RRH in the following position: Tip Tilt Pist Sel...
TelLogLBT672a20100624
waiting for better weather... 20100624_043405: act#173 jumped capsens reading to 1.2m Main.ArmandoRiccardi 24 Jun 2010
TelLogLBT672a20100626
Late start because clouds. Open dome at about 9:15 UT Collimated with Z5= 30.65 and Z6= 79.936 Closed loop short time before 2010 06 26 07:50:20.851562 Rip: FASTD...
TelLogLBT672a20100919UT
Day operations Mounting * Mounted the unit on the telescope * REMAIND: formal request to LBTO about mount the Swagelock of the telescope cooling lines o...
TelLogLBT672a20100920UT
Daytime RRH installation * removed brackets and pins of the mirror cover. Stored in AO cabinet #2 * installed brackets of RRH and installed RRH. No problem...
TelLogLBT672a20100921UT
Day operations Trigger fiber issue * skype con with Norm. Trigger fiber connected in Upper Left Tree House, LIU0602, blade 4, position 2. See IssueTrack#2846 ...
TelLogLBT672a20100922UT
Day time (extended during night because bad weather) Flattening and mode acquisition * daytime for MODS. MODS have control on the status of the telescope. We ...
TelLogLBT672a20100923UT
Day time Flattening XXX Mode acquisition XXX Resume of the mode acquisitions The date part of the tracknumber in the table is "20100922_" TrackNum from t...
TelLogLBT672a20101019UT
"Retracted swing arm" detected by error (false detection) the false detection caused a RIP of the shell adsecarb.R ERR 36764 2010 10 19 01:18:33.449487 MAIN ...
TelLogLBT672a20110116UT
Giornata di test DISASTROSA! Un esempio di probrema con le slope con valori enormi e' visibile, per esempio, nel loop chiuso seguente (vedi idlctrl.R.000012952446...
TelLogLBT672a20110119UT
Astigmatism characterization reproducibility of astigmatism at 90deg after slewing to 26deg * el=90deg * closed loop, bin 2 * saved data: trackin...
TelLogLBT672a20110705a
Mario's communication test on unit LBT672a LBT672a showd problems durin last night, moreover there are hints that communications has problem triggering problems o...
TelLogLBT672a20110817a
Unit A. 17 Aug 2011 activities. * Arrival on the mountain at 1pm (Armando Marco) * checked status of cooling and power (with Micheal and Doug): * p...
TelLogLBT672a20110818
Unit A: 18 Aug 2011 activities * shell set test: all OK Main.ArmandoRiccardi 19 Aug 2011
TelLogLBT672a20110819
Unit A: 19 Aug 2011 activities * no anctivity on unit A today Main.ArmandoRiccardi 19 Aug 2011
TelLogLBT672a20130826
Mon Aug 26 03:58:51 GMT 2013 Start Observations: 10:00 pm Telescope at zenith, HBS running added act 501 in act_wo_icc list. Modification installed. Flattening ...
TelLogLBT672a20130827
Star at 8:00pm Check the trigger in the UTH, the fiber is changed by the spare and the check test is good. The spare fiber is left connected. Original fiber is ch...
TelLogLBT672b20110817
Unit B: 17 Aug activities * Unit powering on: * no fastlink communication: wrong fiber connected to the fastlink port of switch bcu (was unit A configura...
TelLogLBT672b20110818
Unit B: 18 Aug 2011 activities * Hub related work (with Mike Wagner) * move Hub from transportation stand to installation cart * check of RRH inter...
TelLogLBT672b20110829
Unit B: 29 Ago 2011, Telescope Activity Log * Installation of Mirror Unit B on the telescope (SX side) * Connection to Power (from Robert's information) ...
TelLogLBT672b20120408
Lo specchio con il solo set della shell Rippa di continuo senza fare nulla Analisi dei Dump della Master: Dopo il setting, in condizioni statiche: idlctrl.L ...
TelLogLBT672b20120421
su segnalazione di JC trovato specchio in panic con SigGen del crate#1 con il solito problema l'inizio del problema: idlctrl.L INF 139005 2012 04 21...
TelLogLBT672b20120918
analisi di qaulche rip atraverso i dump della master /local/aolog/current/ADAPTIVE SECONDARY_00001347938197_20120918_031637.log durante un CL a 36 modi i modi 10...
TelLogLBT672b20120919
Controllare il RIP: fastdiagn.L ERR 734492 2012 09 19 03:43:28.285363 FUNCTALARM FunctAlarm CHINTCONTROLCURRENT 0354 0.790974 fastdiagn.L ...
TelescopeLogLBT672a
LBT672a related Logs Mar 2010 Log. First Seeing Limited Run * TelLog672a20100311 * TelLogLBT672a20100312 First night * TelLogLBT672a20100312Day * Tel...
TelescopeLogLBT672b
LBT672b related Logs Tables of Opt Int Mat and Flattening Commands Ago/Sep 2011 Log. First Seeing Limited Run * TelLogLBT672b20110829 * TelLogLBT672b201108...
TowerTest20090730
In order to test all the software and hardware that will be used for the influence function measurements with trigger, we started a test session * Installation...
TowerTest20090808
Switched to W1 electronics box, because W2 was continuously resetting the BCU39 board. Mirror configuration NOTE: ICC is active (control current integrator) Remi...
TowerTest20090810
IF acquisition 2009_08_10 Measurement sets to check functionality and configurations preliminary operations * AdSec unit switched on and thermalized * a n...
TowerTest20090812
Karhunen Loève modes optical verification KL matrix produced by Fernando by fitting together zonal mirror modes optical influence functions measurements The KL m...
TowerTest20090817
Hexapod 499.228, 2127.63, 147.621, 159.995, 35.8124, 0.010940 WFS frame rate 25Hz Mirror reflattening Pupil moved with respect to detector mask. Moved pupil...
WhereAreData
Location of data and configurations LBT672 paths adsec_path: HOME STRING '/usr/local/adopt/idl/' DATA STRING '/usr/local/adopt/conf/adsec/672a/data/' CONF STRING...
Number of topics: 177
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback